Can you help me with the following MARKOV CHAINS problem, explaining the solution step by step
Families in a certain country are classified according to whether they reside in rural, urban or suburban areas. Demographic mobility studies estimate that, on average, over the course of a year, 15% of urban families change residence and move to a suburban area, and 5% to a rural area; while 6% of families residing in suburban areas move to urban areas, and 4% to rural areas, and finally 4% of rural families migrate to urban areas and 6% to suburban areas.
a. Define states and transition matrix
b. What is the probability that a family that now lives in an urban area will continue to live in an urban area within two years? And in a suburban? And in a rural one?
c. Suppose that presently 40% of the country's families live in urban areas, 35% in suburban areas, and 25% in rural areas. What percentage of families will live in urban areas in two years?
d. What population distribution should be anticipated in the future if trends do not change?
Answer #1

a. \ States: \ 1-Living\ in\ urban\ area, \ 2-Living\ in\ suburban\ area\\ 3-Living\ in\ rural\ area \\ Transition\ matrix := \ P^{(1)}=\begin{bmatrix} (1\to1) & (1\to 2) & (1\to 3) \\ (2\to 1) &(2\to2) & (2\to3) \\ (3\to1) & (3\to2) & (3\to3) \end{bmatrix} = \begin{bmatrix} 0.8 & 0.15 & 0.05 \\ 0.06 & 0.9 & 0.04 \\ 0.04 & 0.06 & 0.9 \end{bmatrix}

b.\ We\ want\ 2-step\ transition\ matrix\ P^{(2)}\ . \\ P^{(2)}=P^{(1)}P^{(1)} = \begin{bmatrix} 0.6510 &0.2580 & 0.091 \\ 0.1036 & 0.8214 & 0.075 \\ 0.0716 & 0.114 & 0.8144 \end{bmatrix} , \ \ so\ ans : P^{(2)}_{11}=0.6510=urban\ to\ urban\ in\ 2\ years, \\ P^{(2)}_{12}=0.2580=urban\ to\ suburb\ in\ 2\ yrs , \ \ \ P^{(2)}_{13}=0.091 = urban\ to\ rural\\ in\ 2\ yrs.

c.\ We\ know\ current\ distribution\ a^{(0)}=[0.4, \ 0.35, \ 0.25]. We\ want\ distribution\\ after\ 2\ years. \\ a^{(2)} = a^{(0)}P^{(2)}=\begin{bmatrix} 0.4 & 0.35 & 0.25 \end{bmatrix} \begin{bmatrix} 0.6510 &0.2580 & 0.091 \\ 0.1036 & 0.8214 & 0.075 \\ 0.0716 & 0.114 & 0.8144 \end{bmatrix} = [0.31456, \ 0.41919, \ 0.26625] \\ so\ 31.456\%\ of\ families\ will\ live\ in\ urban\ areas\ in\ two\ years.

d. \ let\ future\ steady\ state\ distribution\ be\ \pi_{1}, \pi_{2}, \ \pi_{3} \\ then \ , \ \pi_{1}+\pi_{2}+\pi_{3}= 1 \\ also, \ \pi_{1}=0.8\pi_{1} + 0.06\pi_{2} + 0.04\pi_{3} \\ \pi_{2} = 0.15\pi_{1} + 0.9\pi_{2} + 0.06\pi_{3} \\ solving\ above\ 3\ eqns\ we\ get, \\ \pi_{1} = 0.25969 , \ \pi_{2}=0.36431 , \ \pi_{3}=0.376